OpenStudy (anonymous):

how to solve : x+y=10 25x+50y=.40(10)

OpenStudy (anonymous):

Solve for a variable. Lets use the first statement. x+y=10 x=10-y Now, use this understanding of what "x equals" and apply it to the second statement. x=10-y, therefore 25(10-y)+50y=4 250-25y+50y=4 250+25y=4 25y=-246 y=61.5 Use this numerical value of y in the simplest statement to solve for the other numerical value of x. x+61.5=10 x=(10-61.5) This system is solved for x=(10-61.5) and y=61.5

OpenStudy (anonymous):